# Holographic Tensor Networks as Tessellations of Geometry

### v2

@@ 4 Factorized PEE tensor network
$|\mathcal{T}(\mathbf{z})\rangle=\mathcal{T}_{\mathbf{k}\mathbf{k^{\prime}}}(\mathbf{z})\left(\bigotimes_{\mathcal{C}_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})}|k_{\mathbf{x}}\rangle|k_{\mathbf{x^{\prime}}}\rangle\right),$ (10)
where $\mathbf{k}$ ($\mathbf{k^{\prime}}$) denotes a collection of indices $\{k_{\mathbf{x}}\}$ ($\{k_{\mathbf{x^{\prime}}}\}$). We define the factorized PEE tensor network in the following:
  • Firstly, we set the tensor $\mathcal{T}_{\mathbf{k}\mathbf{k^{\prime}}}(\mathbf{z})$ to be a tensor product of two-index tensors, each of which is associated to a pair of legs lying within the same complete PEE thread, such that the quantum state (8) becomes a product of states associated to each pair of legs:
    $|\mathcal{T}(\mathbf{z})\rangle=\bigotimes_{\mathcal{C}_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})}\mathcal{T}_{ab}(\mathbf{z})|k_{\mathbf{x}}\rangle|k_{\mathbf{x^{\prime}}}\rangle\,.$ (11)
  • Furthermore, we require the state $\mathcal{T}_{ab}(\mathbf{z})|k_{\mathbf{x}}\rangle|k_{\mathbf{x^{\prime}}}\rangle$ to be a pair of maximally entangled qudits, i.e. EPR pairs.
@@
which is also a two-index tensor representing a pair of maximally entangled qudits. The contraction gives us a complete PEE thread with the two open legs tangent to $\mathcal{C}(\mathbf{x},\mathbf{y})$ in the EPR state $|\psi\rangle_{\mathbf{x}\mathbf{y}}=T_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}}$, (see Fig.3). After the contraction along all the PEE threads, we get a tensor network representing the quantum state for all the boundary open legs:
@@
Now we consider any boundary region $A$ in the state (13), and compute the entanglement entropy for $A$. It is convenient to classify all the PEE threads into three classes: (a) $\mathcal{C}^{A}$: threads that anchor only on $A$; (b) $\mathcal{C}^{\bar{A}}$: threads that anchor only on $\bar{A}$; (c) $\mathcal{C}^{A\bar{A}}$: threads that anchor on both $A$ and $\bar{A}$ at each end. Let us rewrite $|\Psi\rangle=|T^{A}\rangle\otimes|T^{\bar{A}}\rangle\otimes|T^{A\bar{A}}\rangle$ where
@@
$=$ $\bigotimes_{\mathbf{x}\in A_{1}}\frac{\mathbbm{1}_{\mathbf{x}}}{v}=\frac{\mathbbm{1}_{A_{1}}}{v^{\mathcal{I}(A,\bar{A})}},$ (18)
where $A_{1}$ denote the qudits in $A$ that are connected to $\bar{A}$ by the threads in $\mathcal{C}^{A\bar{A}}$, and $\mathcal{I}(A,\bar{A})$ counts the number of PEE threads in $\mathcal{C}^{A\bar{A}}$. Therefore, the entanglement entropy between $A$ and $\bar{A}$ is given by:
$S_{A}=S(\rho^{(A\bar{A})}_{A})=\mathcal{I}(A,\bar{A}),$ (19)
where we have set $\log v=1$ such that each thread gives unit contribution to $S_{A}$.
It has been demonstrated that Lin:2024dho , given a static boundary interval $A$ (or any spherical region in higher dimensions) and the corresponding RT surface $\gamma_{A}$, the PEE threads in $\mathcal{C}^{A\bar{A}}$ are the only threads that intersect with $\gamma_{A}$, and each thread only intersects once. This means the number of intersections between $\gamma_{A}$ and the PEE network is exactly $\mathcal{I}(A,\bar{A})$. Since the area of $\gamma_{A}$ can be computed by counting the number of intersections (3), we have
@@ 5 Random PEE tensor network
In the second model, we also start from the PEE network, but set the states $|\mathcal{T}(\mathbf{z})\rangle$ (8) on the bulk vertices to be unit random states following Hayden:2016cfa . We will take the random average of such states when computing the entropy quantities. To be specific, we take an arbitrary reference state $|0_{\mathbf{z}}\rangle$ and define $|\mathcal{T}(\mathbf{z})\rangle=U_{\mathbf{z}}|0_{\mathbf{z}}\rangle$ where $U_{\mathbf{z}}$ is a unitary operator, then the random average is equivalent to an integration over $U_{\mathbf{z}}$ according to the Haar probability measure. Again, after contracting all the internal legs of adjacent bulk sites along all the PEE threads using (9), we obtain the random PEE tensor network representing a boundary state for the boundary open legs.
@@
where $\Sigma_{A}$ is a surface in the network homologous to $A$, and $\mathcal{N}(\Sigma_{A})$ counts the number of contracted tensor legs cut by $\Sigma_{A}$. Note that, the above result applies to general network graphs, including our PEE network. In this case, counting the minimal number of contracted legs cut by $\Sigma_{A}$ is equivalent to computing the area of the minimal homologous surface $\gamma_{A}$, which is just the RT surface. Note that, in order to obtain the result (21), we need to set the contribution $\log v$ to the entropy from each thread to be a large number. Under this setup, we need to divide the density of the PEE threads by an overall constant $\log v$ to keep the boundary PEE unchanged. In other words, the density of intersections between any bulk surface and the PEE network becomes $1/(4G\log v)$, such that $\min\mathcal{N}(\Sigma_{A})=\text{Area}(\gamma_{A})/(4G\log v)$. Hence we reproduced the RT formula,
@@ 6 Discussions
In this work, we assign quantum states to the vertices in the PEE network of AdS space and build two PEE tensor network models, where the entanglement entropy can be explicitly computed. In our models, counting the minimal number of cuts in the tensor network is exactly computing the area of the RT surface in the semi-classical geometry of gravity. This provides the first example of taking natural continuous limit for a discrete tensor network to perfectly match the geometric background.
@@
More interesting models can be developed by including coupling among the PEE threads. In the appendix B, besides regarding the tensors as tensor product of two-index tensors, we also construct a HaPPY-like Pastawski:2015qua PEE tensor network, where we associate a perfect tensor to each site in the PEE network. The RT formula for connected regions in AdS3 is derived using a similar greedy algorithm developed in Pastawski:2015qua for our HaPPY-like PEE tensor network.
As in Hayden:2016cfa ; Pastawski:2015qua , we can also set open legs for the bulk sites to built a holographic code mapping the bulk states to the boundary states, then we can study the bulk quantum correction to the holographic entanglement entropy and identify the analogue of the quantum extremal surface Engelhardt:2014gca . Perhaps the most interesting future direction is to generalize the PEE tensor network to geometric background beyond the Poincaré AdS space, thus construct a new framework to study holographic correspondence beyond AdS/CFT. This is promising because the Crofton formula works for more generic Riemann manifolds, and we can build the network of all the geodesics in the Kinematic Space. The generalization will be non-trivial, as it is quite possible that the geodesics in the bulk will not always anchor on the boundary, which may correspond to spaces containing black holes or singularities.
@@ Appendix A The RT formula in the factorized PEE tensor network
Before making a comparison between the factorized PEE tensor network with the HaPPY code Pastawski:2015qua , we first introduce some preliminary about isometric tensor for the section to be self-contained.
@@
It is also useful to derive the RT formula using the arguments in Pastawski:2015qua for more generic tensor network models. To be specific, let us cut the tensor network open along a homologous surface $\mathcal{C}$ whose boundary matches the boundary of $A$, such that the tensor network is divided into two parts $\mathcal{P}$ and $\mathcal{Q}$, which are maps from $\mathcal{C}$ to $A$ and from $\mathcal{C}$ to $\bar{A}$ respectively. The boundary state can be written as
[…]
whose corresponding entanglement entropy $S_{A}$ is upper bounded by the dimension of the Hilbert space on $\mathcal{C}$, i.e. the number of cuts $\mathcal{N}(\mathcal{C})$ between $\mathcal{C}$ and the network. Then we get the minimal upper bound by choosing the homologous surface with minimal number of cuts, which we also denote as $\gamma_{A}$, hence
$S_{A}\leq\mathcal{N}(\gamma_{A})\log v.$ (32)
Furthermore, according to the proposition (A.3), if the tensors $\mathcal{P}$ and $\mathcal{Q}$ are isometries, then $|\mathcal{Q}_{\mathbf{c}}\rangle$ and $|\mathcal{P}_{\mathbf{c}}\rangle$ are two sets of orthonormal vectors. This implies $\rho_{A}$ is proportional to the identity such that the above inequality is saturated.
@@
Appendix B HaPPY-like PEE tensor network
Besides regarding the tensors as tensor product of two-index tensors, we can also construct a HaPPY-like PEE tensor network, where we associate a perfect tensor on each site in the PEE network. Firstly, let us recall the definition of perfect tensor as follows:
Definition B.1.
As was shown in Pastawski:2015qua , a bulk tensor network built from perfect tensors also satisfy the entropy bound (32), and the inequality is saturated if and only if the two tensors $\mathcal{P}$ and $\mathcal{Q}$ obtained by cutting the whole bulk tensor along $\gamma_{A}$ are isometric. Since the graph of our network is the PEE network, again we have $\mathcal{N}(\gamma_{A})={\text{Area}(\gamma_{A})}/{4G}$. In order to reproduce the RT formula, we only need to demonstrate that the mapping $\mathcal{P}$ is an isometry that maps $\gamma_{A}$ to $A$ and $\mathcal{Q}$ is an isometry that map $\gamma_{A}$ to $\bar{A}$.
Now we build the isometric map from $\gamma_{A}$ to $A$ for our HaPPY-like PEE tensor network using an analogous greedy algorithm proposed in Pastawski:2015qua . The procedure is decomposed into the following steps:
  1. 1.
    Given any connected boundary region $A$, let us consider a homologous surface $\Sigma_{A}$ which cuts the PEE tensor network open, and we denote the region enclosed by $A$ and $\Sigma_{A}$ as $\mathcal{W}_{\Sigma_{A}}$ (we do not include the sites on $\Sigma_{A}$ into $\mathcal{W}_{\Sigma_{A}}$). In Fig.5, we mark $\mathcal{W}_{\Sigma_{A}}$ as the blue region and $\Sigma_{A}$ as the dashed red line. Note that, we require the tensor in $\mathcal{W}_{\Sigma_{A}}$ to be an isometric map from the legs on $\Sigma_{A}$ to those on $A$. More explicitly, here the legs on $\Sigma_{A}$ means the legs emanating from $\Sigma_{A}$ into the blue region.
  2. 2.
  3. 3.
    If the two dashed legs also stay in $\mathcal{W}_{\Sigma_{A}}$, then the number of legs in $\mathcal{W}_{\Sigma_{A}}$ is greater than the number of outside legs (see Fig.5a), hence the perfect tensor on this site represents an isometric map from the outside legs to the inside legs. In this case we extend $\mathcal{W}_{\Sigma_{A}}$ to swallow this site, hence the extended region still corresponds to a isometric map from the new $\Sigma_{A}$ to $A$.
  4. 4.
  5. 5.
  6. 6.

### v3

@@ 4 Factorized PEE tensor network
$|\mathcal{T}(\mathbf{z})\rangle=\mathcal{T}_{\mathbf{a}\mathbf{b}}(\mathbf{z})\left(\bigotimes_{\mathcal{C}_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})}|a_{\mathbf{x}}\rangle_{\mathbf{z}}|b_{\mathbf{x^{\prime}}}\rangle_{\mathbf{z}}\right),$ (10)
where $\mathbf{a}$ ($\mathbf{b}$) denotes a collection of indices $\{a_{\mathbf{x}}\}$ ($\{b_{\mathbf{x^{\prime}}}\}$). We define the factorized PEE tensor network in the following:
  • Firstly, we set the tensor $\mathcal{T}_{\mathbf{a}\mathbf{b}}(\mathbf{z})$ to be a tensor product of two-index tensors, each of which is associated to a pair of legs lying within the same complete PEE thread, such that the quantum state (8) becomes a product of states associated to each pair of legs:
    $|\mathcal{T}(\mathbf{z})\rangle=\bigotimes_{\mathcal{C}_{\mathbf{z}}(\mathbf{x},\mathbf{x^{\prime}})}\mathcal{T}_{ab}(\mathbf{z})|a_{\mathbf{x}}\rangle_{\mathbf{z}}|b_{\mathbf{x^{\prime}}}\rangle_{\mathbf{z}}\,.$ (11)
  • Furthermore, we require the state $\mathcal{T}_{ab}(\mathbf{z})|a_{\mathbf{x}}\rangle_{\mathbf{z}}|b_{\mathbf{x^{\prime}}}\rangle_{\mathbf{z}}$ to be a pair of maximally entangled qudits, i.e. EPR pairs.
@@
which is also a two-index tensor representing a pair of maximally entangled qudits. The contraction gives us a complete PEE thread with the two open legs tangent to $\mathcal{C}(\mathbf{x},\mathbf{y})$ in the EPR state $|\psi\rangle_{\mathbf{x}\mathbf{y}}=T_{ab}(\mathbf{x},\mathbf{y})|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}}$, (see Fig.3). Note that, since $\mathbf{x}$ and $\mathbf{y}$ are boundary sites, we cannot use them to parameterize their legs since by default there are numerous qudits with open leg lying in the same boundary site. Just keep in mind that, here $|a\rangle_{\mathbf{x}}|b\rangle_{\mathbf{y}}$ only denotes the state of the pair of open legs that is tangent to the corresponding PEE thread $\mathcal{C}(\mathbf{x},\mathbf{y})$ respectively. After the contraction along all the PEE threads, we get a tensor network representing the quantum state for all the boundary open legs:
@@
Now we consider any boundary region $A$ in the state (13), and compute the entanglement entropy for $A$. It is convenient to classify all the PEE threads into three classes: (a) $\mathcal{C}^{A}$: threads that anchor only on $A$; (b) $\mathcal{C}^{\bar{A}}$: threads that anchor only on $\bar{A}$; (c) $\mathcal{C}^{A\bar{A}}$: threads that anchor on both $A$ and $\bar{A}$ at each end. See the Fig.5a for illustration. Let us rewrite $|\Psi\rangle=|T^{A}\rangle\otimes|T^{\bar{A}}\rangle\otimes|T^{A\bar{A}}\rangle$ where
@@
$=$ $\bigotimes_{\mathcal{C}(\mathbf{x},\mathbf{y})\in\mathcal{C}^{A\bar{A}}}\frac{\mathbbm{1}_{\mathbf{x}}}{v}.$ (18)
In order to match the case of AdS/CFT, the total flux of PEE threads connecting $A$ and $\bar{A}$ should match to $\mathcal{I}(A,\bar{A})$. Again, let us set $\log v=1$, then the number of PEE threads in $\mathcal{C}^{A\bar{A}}$ coincides with $\mathcal{I}(A,\bar{A})$, and the entanglement entropy between $A$ and $\bar{A}$ is given by:
$S_{A}=S(\rho^{(A\bar{A})}_{A})=\mathcal{I}(A,\bar{A}).$ (19)
It has been demonstrated that Lin:2024dho , given a static boundary interval $A$ (or any spherical region in higher dimensions) and the corresponding RT surface $\gamma_{A}$, the PEE threads in $\mathcal{C}^{A\bar{A}}$ are the only threads that intersect with $\gamma_{A}$, and each thread only intersects once. This means the number of intersections between $\gamma_{A}$ and the PEE network is exactly $\mathcal{I}(A,\bar{A})$. Furthermore, since the area of $\gamma_{A}$ is computed by counting the number of intersections (3), we have
@@
5 HaPPY-like PEE tensor network
Besides regarding the tensors as tensor product of two-index tensors, we can also construct a HaPPY-like PEE tensor network following Pastawski:2015qua , where we associate a perfect tensor on each site in the PEE network. Firstly, let us recall the definition of perfect tensor as follows:
Definition 5.1.
Let us start from a generic tensor network with no open legs in the bulk, and consider a connected boundary region $A$. Furthermore, let us cut the tensor network open along a homologous surface $\mathcal{C}$ whose boundary matches the boundary of $A$, such that the tensor network is divided into two parts $\mathcal{P}$ and $\mathcal{Q}$, which are maps from $\mathcal{C}$ to $A$ and from $\mathcal{C}$ to $\bar{A}$ respectively. The boundary state can be written as
[…]
whose corresponding entanglement entropy $S_{A}$ is upper bounded by the log of the Hilbert space dimension of $\mathcal{C}$, i.e. $\mathcal{N}(\mathcal{C})\log v$ where $\mathcal{N}(\mathcal{C})$ is the number of intersections between $\mathcal{C}$ and the network. Then we get the minimal upper bound by choosing the homologous surface with minimal number of cuts, which we also denote as $\gamma_{A}$, hence
$S_{A}\leq\mathcal{N}(\gamma_{A})\log v.$ (23)
As demonstrated in Pastawski:2015qua , the above inequality is saturated if and only if the two tensors $\mathcal{P}$ and $\mathcal{Q}$, obtained by cutting the entire bulk tensor along $\gamma_{A}$, are isometric. We now apply this discussion to our PEE network constructed from perfect tensors, and again set $\log v=1$ so that $\mathcal{N}(\gamma_{A})=\text{Area}(\gamma_{A})/{4G}$. To reproduce the RT formula, it suffices to show that the mapping $\mathcal{P}$ is an isometry from $\gamma_{A}$ to $A$, and $\mathcal{Q}$ is an isometry from $\gamma_{A}$ to $\bar{A}$.
Now we prove the isometric property of $\mathcal{P}$ and $\mathcal{Q}$ using an analogous greedy algorithm proposed in Pastawski:2015qua . The procedure is decomposed into the following steps:
  1. 1.
    Given any connected boundary region $A$, let us consider a homologous surface $\Sigma_{A}$ which cuts the PEE tensor network open, and we denote the region enclosed by $A$ and $\Sigma_{A}$ as $\mathcal{W}_{\Sigma_{A}}$ (we do not include the sites on $\Sigma_{A}$ into $\mathcal{W}_{\Sigma_{A}}$). In Fig.4, we mark $\mathcal{W}_{\Sigma_{A}}$ as the blue region and $\Sigma_{A}$ as the dashed red line. Note that, we start from a $\mathcal{W}_{\Sigma_{A}}$ that is an isometric map from the legs on $\Sigma_{A}$ to those on $A$. Here, the legs on $\Sigma_{A}$ means the legs emanating from $\Sigma_{A}$ into the blue region.
  2. 2.
  3. 3.
    If the two dashed legs also stay in $\mathcal{W}_{\Sigma_{A}}$, then the number of legs in $\mathcal{W}_{\Sigma_{A}}$ is greater than the number of outside legs (see Fig.4a), hence the perfect tensor on this site represents an isometric map from the outside legs to the inside legs. Since the composition of two isometric maps is also isometric, we extend $\mathcal{W}_{\Sigma_{A}}$ to incorporate this site. Consequently, the extended region still corresponds to an isometric map from the new boundary $\Sigma_{A}$ to $A$.
  4. 4.
  5. 5.
  6. 6.
@@ 6 Random PEE tensor network
In the second model, we also start from the PEE network, but set the states $|\mathcal{T}(\mathbf{z})\rangle$ (8) on the bulk vertices to be unit random states following Hayden:2016cfa . We will take the random average of such states when computing the entropy quantities. To be specific, we take an arbitrary reference state $|0\rangle_{\mathbf{z}}$ and define $|\mathcal{T}(\mathbf{z})\rangle=U_{\mathbf{z}}|0\rangle_{\mathbf{z}}$ where $U_{\mathbf{z}}$ is a unitary operator, then the random average is equivalent to an integration over $U_{\mathbf{z}}$ according to the Haar probability measure. Again, after contracting all the internal legs of adjacent bulk sites along all the PEE threads using (9), we obtain the random PEE tensor network representing a boundary state for the boundary open legs.
@@
where $\Sigma_{A}$ is a surface in the network homologous to $A$, and $\mathcal{N}(\Sigma_{A})$ counts the number of contracted tensor legs cut by $\Sigma_{A}$. Note that, the above result applies to general network graphs, including our PEE network. In this case, counting the minimal number of contracted legs cut by $\Sigma_{A}$ is equivalent to computing the area of the minimal homologous surface $\gamma_{A}$, which is just the RT surface.
Note that, in order to obtain the result (24), we need to set the contribution $\log v$ to the entropy from each thread to be a large number. Under this setup, we need to divide the density of the PEE threads by an overall constant $\log v$, such that the boundary PEE structure and the total crossing of the bulk surfaces are unchanged. In other words, the density of intersections between any bulk surface and the PEE network becomes $1/(4G\log v)$, such that $\min\mathcal{N}(\Sigma_{A})=\text{Area}(\gamma_{A})/(4G\log v)$. Hence we reproduced the RT formula,
@@ 7 Discussions
In this work, we assign quantum states to the vertices in the PEE network of AdS space and build three PEE tensor network models, where the entanglement entropy can be explicitly computed. In our models, counting the minimal number of cuts in the tensor network is exactly computing the area of the RT surface in the semi-classical geometry of gravity. This provides the first example of taking natural continuous limit for a discrete tensor network to perfectly match the geometric background. In our construction, the PEE network is determined by the PEE structure of the boundary CFT state; equivalently, it is built from the geodesics in the kinematic space, which itself is an intrinsic property of the background geometry. Thus, the network is not manually adjusted, and the RT formula is not used as an input.
@@
More interesting models can be developed by including coupling among the PEE threads. In section 5, besides regarding the tensors as tensor product of two-index tensors, we also construct a HaPPY-like Pastawski:2015qua PEE tensor network, where we associate a perfect tensor to each site in the PEE network. The RT formula for connected regions in AdS3 is derived using a similar greedy algorithm developed in Pastawski:2015qua for our HaPPY-like PEE tensor network.
As in Hayden:2016cfa ; Pastawski:2015qua , we can also set open legs for the bulk sites to built a holographic code mapping the bulk states to the boundary states, then we can study the bulk quantum correction to the holographic entanglement entropy and identify the analogue of the quantum extremal surface Engelhardt:2014gca . Perhaps the most interesting future direction is to generalize the PEE tensor network to geometric background beyond the Poincaré AdS space, thus construct a new framework to study holographic correspondence beyond AdS/CFT. This is nontrivial, as the Crofton formula requires essential modifications to apply to more general Riemannian manifolds. In such configurations, besides geodesics anchored on the boundary, one must also consider those that terminate at black hole singularities or end in the bulk.
Our work may provide a new framework for studying quantum field theories on a smooth geometric background using continuous tensor network models, without losing the information of background geometry. This could be an advantage compared with the other continuous tensor network models (see also Quijandria:2014gsa ; Tagliacozzo:2014bta ; Draxler:2016nxp ; Ganahl:2016mjv ; DelasCuevas:2017yxc ; Ganahl:2018jma ; Cotler:2018ufx ; Osborne:2019bsq ; Tang:2020vyv ; Tuybens:2020uth ; Karanikolaou:2020rld ; Stottmeister:2020ezd ; Zou:2019xbi ; Witteveen:2019lsk ; Campos:2021zce ; Osborne:2021ppp ; Lukin:2022rds ; Martyn:2022oll ; Schmoll:2023eez ; Vardian:2024jll ; Chemissany:2025vye for an incomplete list of recent developments). It is particularly interesting to apply this framework in flat space, where the Crofton formula holds and most non-gravitational field theories are studied.
@@ Appendix A The RT formula in the factorized PEE tensor network
We first introduce some preliminary about isometric tensor for the section to be self-contained.
@@
It is also useful to derive the RT formula using the arguments in Pastawski:2015qua for more generic tensor network models. Again, let us cut the tensor network open along a homologous surface $\mathcal{C}$ whose boundary matches the boundary of $A$, such that the tensor network is divided into two parts $\mathcal{P}$ and $\mathcal{Q}$, which are maps from $\mathcal{C}$ to $A$ and from $\mathcal{C}$ to $\bar{A}$ respectively. According to the proposition (A.3), if the tensors $\mathcal{P}$ and $\mathcal{Q}$ are isometries, then $|\mathcal{Q}_{\mathbf{c}}\rangle$ and $|\mathcal{P}_{\mathbf{c}}\rangle$ are two sets of orthonormal vectors. This implies $\rho_{A}$ is proportional to the identity such that the upper bound (23) for entanglement entropy is saturated.